# Unveiling the Nature and Fate of the Almost-Dark Cloud AGC 226178 through Mapping

### v1

@@ 6 Acknowledgements
To further explore the distribution across the velocity space, we extract position-velocity (PV) diagrams along five different directions indicated in the bottom-middle panel of Fig. 5, the PV diagrams are shown in Fig. A1. The extraction slits of the PV diagrams are chosen to be either along or parallel to the major and minor axes, where the major axis (slit D) matches the direction of the maximum velocity gradient whereas the minor axis (slit A) goes across the brightest peak. An apparent rotation-like velocity gradient is limited to the major axis. Moreover, the gradient appears to be driven by the two brightest clumps, with faint but notable gas in the ‘forbidden’ quadrants for a rotation-dominated PV diagram.

### v2

@@ Appendix A PV diagram of VLA observation for AGC 226178
To further explore the distribution across the velocity space, we extract position-velocity (PV) diagrams along five different directions indicated in the bottom middle panel of Fig. 5, the PV diagrams are shown in Fig. 14. The extraction slits of the PV diagrams are chosen to be either along or parallel to the major and minor axes, where the major axis (slit D) matches the direction of the maximum velocity gradient whereas the minor axis (slit A) goes across the brightest peak. An apparent rotation-like velocity gradient is limited to the major axis. Moreover, the gradient appears to be driven by the two brightest clumps, with faint but notable gas in the ‘forbidden’ quadrants for a rotation-dominated PV diagram.
Appendix B Properties of clumps in AGC 226178.
| ID | (deg) | (deg) | (km s-1) | (mJy km s-1) | (arcsecarcsec) | (km s-1) | |
| 1 | 191.67907 | 10.36797 | 1581.2 | 58.7 | 12.818.1 | 7.2 | BC3s9, BC3s10, BC3s12, BC3s13, BC3s14 |
| 2 | 191.67497 | 10.36574 | 1591.1 | 56.9 | 16.918.1 | 9.0 | BC3s15, BC3s18, BC3s23 |
| 3 | 191.67847 | 10.37008 | 1570.1 | 29.7 | 16.412.9 | 6.2 | – |
| 4 | 191.67606 | 10.37303 | 1577.3 | 18.8 | 10.811.3 | 5.0 | BC3s4, BC3s24W, BC3s26W |
| 5 | 191.68145 | 10.36332 | 1573.7 | 12.7 | 7.410.1 | 4.0 | BC3s19 |
| 6 | 191.67465 | 10.36463 | 1578.1 | 13.5 | 8.111.9 | 4.0 | BC3s20 |
| 7 | 191.67906 | 10.36520 | 1568.9 | 18.1 | 14.310.9 | 3.2 | – |
| 8 | 191.68031 | 10.36443 | 1591.1 | 4.4 | 5.75.9 | 0.3 | – |
| 9 | 191.67911 | 10.37473 | 1574.3 | 8.1 | 9.58.3 | 2.6 | – |
| 10 | 191.68112 | 10.36895 | 1589.5 | 8.3 | 6.411.2 | 1.5 | – |